OpenStudy (anonymous):

which expression is equivalent to (cos x)(tan(-x)) -sin x sin x -csc x csc x

OpenStudy (anonymous):

\[y=\cos(x)\tan(-x)\]\[\tan(-x)=-\tan(x)\]\[y=-\cos(x)\tan(x)\]\[\tan(x)=\frac{\sin(x)}{\cos(x)}\]\[y=-\cos(x) \times \frac{\sin(x)}{\cos(x)}=??\]